# Large Language Models for Blockchain Security: A Systematic Literature Review

### v3

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools in various domains involving blockchain security (BS). Several recent studies are exploring LLMs applied to BS. However, there remains a gap in our understanding regarding the full scope of applications, impacts, and potential constraints of LLMs on blockchain security. To fill this gap, we conduct a literature review on LLM4BS.
As the first review of LLM’s application on blockchain security, our study aims to comprehensively analyze existing research and elucidate how LLMs contribute to enhancing the security of blockchain systems. Through a thorough examination of scholarly works, we delve into the integration of LLMs into various aspects of blockchain security. We explore the mechanisms through which LLMs can bolster blockchain security, including their applications in smart contract auditing, transaction anomaly detection, vulnerable repair, program analysis of smart contracts, and cryptocurrency community operations. Furthermore, we critically assess the challenges and limitations associated with leveraging LLMs for blockchain security, considering factors such as scalability, privacy concerns, and ethical concerns. Our review sheds light on the opportunities and potential risks of LLM4BS task, providing valuable insights for researchers, practitioners, and policymakers alike.
@@
Blockchain Security , Large Language Model , Survey
PACS:
0000 , $11$
MSC:
0000 , $11$
journal: Infomation Fusion
@@ 1 Introduction
As the digital epoch progresses, the confluence of artificial intelligence with blockchain technology emerges as a groundbreaking development, particularly at the juncture where Large Language Models (LLMs) [1, 2, 3, 4, 5, 6] intersect with the ever-evolving domain of blockchain security [7, 8, 5, 6, 9, 10]. LLMs have catapulted to the forefront of blockchain security [11, 12], demonstrating profound capabilities in the text (esp. source code [13, 14, 6, 15]) generation and comprehension—abilities that mirror human-like proficiency [16, 17]. This transformative impact is attributable to their expansive datasets, sophisticated architectures, and the deep neural networks that underpin their operational frameworks [1, 18, 17].
The robustness of LLMs in discerning and synthesizing complex patterns within data positions them as invaluable assets in enhancing the security measures within blockchain systems [19, 20, 3, 4]. Concretely, the granular analysis of smart contracts [13], the meticulous scrutiny of transactions [21], and Automatically generating code [22] are among the critical tasks that LLMs are adept at performing with remarkable efficacy [14, 23, 9].
However, the path to integrating these cognitive powerhouses into blockchain security is met with an array of challenges that beckon for consideration. From navigating the intricate dynamics of ever-advancing cybersecurity threats to addressing the ethical concerns that accompany AI deployment, the trajectory is as demanding as it is promising. There is still a lack of work to comprehensively depict the current application status and future development prospects of LLM in blockchain security (BS).
Table 1: Table of LLM4BS researches
@@
Smart Contract Auditing 11
SMARTINV [24], GPTScan [13], David et al. [25], Karanjai et al. [26],
ContractArmor [27], Ortu et al. [28], ASSBert [29], PSCVFinder [30],
LLM4Vuln [31], TrustLLMf [32], AuditGPT [33]
@@
Smart Contract Development 5
Storhaug et al. [22], karanjai et al. [26], MazzumaGPT [37], Du et al. [38],
GPTutor [39]
@@
compilers [45], zero-knowledge proofs [46], model training [20, 47],
NFT generation [48]
To fill the gap, our work seeks to delve into the multifaceted role of LLMs within the realm of blockchain security, exploring the comprehensive spectrum of LLM on blockchain security (LLM4BS) tasks. We commence by delineating the contemporary landscape of Large Language Model (LLM) applications across diverse domains §2.1, as well as the myriad of security threats implicated by the blockchain technology §2.2. Then, as we display in Table.1, we elaborate on the incorporation and progression of LLM4BS task, involving smart contract auditing, block transaction detection, contract dynamic analysis, smart contract development, and cryptocurrency community contributors §3. Thereafter, we meticulously select three quintessential cases of LLM4BS tasks to elucidate the state-of-the-art LLM4BS tasks §4, comprising LLM4FUZZ [35], SMARTINV [24], BLOCKGPT [21]. Finally, we present an insightful discourse on the challenges presently faced within the ambit of LLM4BS, and proffer prospective trajectories for future research and development in this emergent field §5.
@@
  • To the best of our knowledge, and after a meticulous review of the extant literature, we can confidently assert that our work represents the first systematic examination focusing on the application of Large Language Models (LLM4BS) to tasks within the realm of blockchain security, offering a pioneering exploration of the interplay between advanced AI and cryptographic ledger systems.
  • In our comprehensive survey, we meticulously chronicle the current landscape of Large Language Model (LLM) applications in the domain of blockchain security, delving into a detailed analysis of how LLMs are employed across various scenarios, from the enhancement of smart contract reliability to the fortification of distributed ledger integrity, thereby shedding light on the multifaceted contributions of this cutting-edge technology.
  • Emanating from our research, we rigorously compile and summarize a suite of practical academic accomplishments pertaining to the application of Large Language Models (LLMs) in fortifying blockchain security, in conjunction with proposing an array of promising and prospective avenues for future research, which we anticipate will catalyze substantial advancements and innovations within this burgeoning intersection of fields.
@@ 2 Overview of LLM4BS
We will provide some basic knowledge about LLM4BS tasks in this section, including LLM applications §2.1 and threats of blockchain security §2.2.
@@ 2.1.1 Definition and Characteristics of LLMs
Furthermore, LLMs demonstrate an advanced understanding of context within language, enabling them to generate responses or predictions that are sensitive to the surrounding textual context [54]. This contextual awareness is achieved through mechanisms such as attention mechanisms and positional encodings, which enable LLMs to attend to relevant parts of the input sequence and model long-range dependencies effectively [55].
@@ 2.1.2 Applications of LLMs in Various Domains
As we display in Fig. 1, the versatility and efficacy of LLMs have led to their widespread adoption across diverse domains and applications, where they have demonstrated exceptional performance and utility [57]. Some notable applications of LLMs include:
@@ 2.2.2 Common Security Threats in Blockchain Systems
Auxiliary Service Vulnerabilities: Auxiliary services, such as wallets, exchanges, oracles, and decentralized applications (DApps), serve as entry points for attackers to exploit vulnerabilities and compromise the security of blockchain systems [104, 105]. Security breaches in auxiliary services, such as exchange hacks, wallet vulnerabilities, or oracle manipulation attacks, can lead to the loss of funds, unauthorized access to user data, or manipulation of on-chain transactions [106, 107]. Furthermore, the interconnected nature of auxiliary services within the blockchain ecosystem amplifies the impact of security breaches, as vulnerabilities in one service can propagate to others, resulting in widespread disruption and financial losse.
@@
3 Taxonomy of LLM4BS task
@@ 3.1 LLM as Code auditor on Smart Contracts
The application of LLM in the domain of smart contract code auditing and vulnerability detection can be succinctly encapsulated as follows: Advanced tools, such as SMARTINV [24], GPTScan [13], David et al. [25], Karanjai et al. [26], ContractArmor [27], Ortu et al. [28], ASSBert [29], PSCVFinder [30], LLM4Vuln [31], TrustLLMf [32] and AuditGPT [33]. The tools powered by Large Language Models in Table.2 signify a monumental shift from traditional, pattern-based analysis methodologies towards more contextually aware and comprehensive inspection techniques. These cutting-edge tools extend their analytical prowess beyond static patterns by knitting together disparate threads of information, including the nuanced aspects of natural language documentation that detail the intended functions and transactional constructs of smart contracts.
This marriage of code and contextual data through a multimodal lens equips such tools with the capacity to unravel complex logical oversights and identify subtle ”machine un-auditable” bugs, which would otherwise evade detection. By assimilating and interpreting the richer tapestry of human language explanations paired with code, LLM-based tools delve deeper into the intricate web of smart contract interactions. The profound understanding garnered from this approach not only sheds light on hidden vulnerabilities but also fortifies smart contracts against the myriad of risks that could lead to substantial financial repercussions.
In essence, the integration of Large Language Models in smart contract analysis marks a significant leap in safeguarding blockchain technology’s infrastructural integrity. It underscores an evolving landscape where artificial intelligence converges with software development practices to bolster security measures. This proactive identification and remediation of weaknesses within smart contracts, facilitated by the keen insights offered by LLMs, are instrumental in cementing trust and reliability in blockchain transactions—hence mitigating potential financial liabilities and reinforcing the bedrock of digital contracts.
@@ 3.2 LLM as Analyzers for abnormal transaction
The application of LLMs for blockchain transaction analysis, such as BLOCKGPT [21] and Nicholls et al. [34], underscores their crucial role in conducting real-time monitoring to detect signs of irregular or suspicious behavior. These tools in Table 3 represents a significant advancement in the field, as they provide a more dynamic and adaptable approach to identifying potential threats within blockchain transactions.
@@
In addition to identifying potentially fraudulent transactions, LLMs also contribute to risk assessment and regulatory compliance. By analyzing the transaction data against current compliance standards and risk models, LLMs can assist financial institutions in managing their risk exposure and adhering to anti-money laundering (AML) and know your customer (KYC) regulations. Their sophisticated analysis capabilities can provide valuable insights to compliance officers and regulatory bodies, allowing for a more proactive approach to detecting and preventing financial crimes.
@@ 3.3 LLM as Fuzzer for Smart Contract
Large Language Models (LLMs) have been increasingly employed to elevate the process of fuzzing, particularly in the realm of smart contract security analysis ,such as LLM4FUZZ [35] and ACFIX [36]. This methodology in Table 4 involves utilizing LLMs to accurately assess the complexity and vulnerability likelihood of specific code regions within a smart contract. Consequently, these metrics serve to guide the direction and focus of fuzzers, steering them towards code segments that are more likely to harbor potential security threats.
@@
Moreover, this refined fuzzing technique allows for the integration of user-defined invariants—manually inserted assertions to monitor and manage the state during fuzzing. By doing so, it reduces the exploration overhead and improves the detection of more profound logical issues that regular fuzzing routines might miss. Evaluations of this LLM-enhanced fuzzing method within real-world decentralized finance (DeFi) projects have demonstrated its effectiveness, outperforming baseline fuzzing parameters and uncovering significant vulnerabilities. These vulnerabilities, if left undetected and exploited, could potentially result in substantial financial losses.
@@ 3.4 LLM as Developer for Smart Contract
Storhaug et al. [22] Vulnerability-constrained decoding approach GPT-J-6B ISSRE,2023
@@
Recent studies in Table 5, such as Storhaug et al. [22], karanjai et al. [26], MazzumaGPT [37], Du et al. [38] and GPTutor [39], have begun to scrutinize the efficacy and reliability of Large Language Models (LLMs) like ChatGPT and Google Palm2 in the automated generation of smart contracts. These smart contracts are integral to the blockchain ecosystem, executing agreements without the need for intermediaries, and their accuracy and security are paramount. The research primarily constructs a testing framework that assesses smart contracts on multiple fronts — validity, correctness, efficiency, security, and maintainability.
@@
These works point to the necessity for comprehensive analysis and improvement in the methodologies employed by LLMs. There is optimism that future iterations of LLMs, with better training and prompt-design considerations, could enhance the quality and security of AI-generated smart contracts. It also hints at the potential for these tools to revolutionize contract generation by reducing the time and effort required, while flagging the urgent need for more robust security measures and testing methods.
@@ 3.5 LLM as participants for Cryptocurrency community
Governance emerges as a major theme, as LLMs could contribute significantly to the structuring and transparency of this largely unregulated space. The first document outlines the broader governance challenges faced by AI systems, suggesting blockchain as a viable solution to introduce verifiability and accountability. On the other hand, limitations of LLMs in capturing the complexities of legal reasoning are highlighted, a concern that is echoed across the three studies to varying degrees.
@@
In essence, the collective narrative from the three documents converges on the premise that LLMs hold transformative potential for the cryptocurrency community’s governance and legal sectors but must overcome challenges in understanding and application before they can be fully trusted in autonomous roles.

### v4

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools across various domains within cyber security. Notably, recent studies are increasingly exploring LLMs applied to the context of blockchain security (BS). However, there remains a gap in a comprehensive understanding regarding the full scope of applications, impacts, and potential constraints of LLMs on blockchain security. To fill this gap, we undertake a literature review focusing on the studies that apply LLMs in blockchain security (LLM4BS).
Our study aims to comprehensively analyze and understand existing research, and elucidate how LLMs contribute to enhancing the security of blockchain systems. Through a thorough examination of existing literature, we delve into the integration of LLMs into various aspects of blockchain security. We explore the mechanisms through which LLMs can bolster blockchain security, including their applications in smart contract auditing, transaction anomaly detection, vulnerability repair, program analysis of smart contracts, and serving as participants in the cryptocurrency community. Furthermore, we assess the challenges and limitations associated with leveraging LLMs for enhancing blockchain security, considering factors such as scalability, privacy concerns, and ethical concerns. Our thorough review sheds light on the opportunities and potential risks of tasks on LLM4BS, providing valuable insights for researchers, practitioners, and policymakers alike.
@@
Blockchain Security , Large Language Model , Literature Review
@@ 1 Introduction
As the digital era advances, the confluence of artificial intelligence with blockchain technology emerges as a groundbreaking development, particularly at the juncture where Large Language Models (LLMs) [1, 2, 3, 4, 5, 6] intersect with the ever-evolving domain of blockchain security [7, 8, 5, 6, 9, 10]. LLMs have risen to the forefront of blockchain security [11, 12], showcasing profound capabilities in text generation and comprehension [13, 14, 6, 15]), especially in source code analysis. These abilities mirror human-like proficiency [16, 17]. This transformative impact is attributable to their expansive datasets, sophisticated architectures, and the deep neural networks that underpin their operational frameworks [1, 18, 17].
The robustness of LLMs in discerning and synthesizing complex patterns within data positions them as invaluable assets in enhancing the security measures within blockchain systems [19, 20, 3, 4, 21, 22, 23, 24, 25]. Concretely, the granular analysis of smart contracts [13], the meticulous scrutiny of transactions [26], and automatic code (resp. text) generation [27] are among the critical tasks that LLMs are adept at performing with remarkable efficacy [14, 28, 9].
However, integrating these cognitive powerhouses into blockchain security is met with an array of challenges that beckon for consideration. Navigating the intricate dynamics of ever-advancing cybersecurity threats and addressing the ethical concerns that accompany AI deployment make this trajectory as demanding as it is promising. Despite the progress, there is still a lack of comprehensive work depicting the current application status and future development prospects of LLM in blockchain security (BS).
Table 1: Table of LLM4BS studies
@@
Smart Contract Auditing 15
SMARTINV [29], GPTScan [13], David et al. [30],
Karanjai et al. [31], ContractArmor [32], Ortu et al. [33],
ASSBert [34], PSCVFinder [35], LLM4Vuln [36],
TrustLLMf [37], AuditGPT [38], PropertyGPT [39]
Chen et al. [40], Jain et al. [41] and SolGPT [42]
@@
Smart Contract Development 8
Storhaug et al. [27], karanjai et al. [31], MazzumaGPT [46],
Du et al. [47], GPTutor [48], Petrovic et al. [49],
Zhao et al. [50] and Haque et al. [51]
@@
compilers [57], zero-knowledge proofs [58],
model training [20, 59] and NFT generation [60]
To fill the gap, we seek to delve into the multifaceted role of LLMs within the realm of blockchain security, exploring the comprehensive spectrum of LLMs on blockchain security (LLM4BS) tasks. We commence by delineating the contemporary landscape of Large Language Model (LLM) applications across diverse domains (§2.1), as well as the myriad of security threats implicated by the blockchain technology (§2.2). Then, as illustrated in Table.1, we elaborate on the incorporation and progression of LLM4BS tasks, involving smart contract auditing, block transaction detection, contract dynamic analysis, smart contract development, and cryptocurrency community contributors (§3). Thereafter, we meticulously select three quintessential cases of LLM4BS tasks to elucidate the state-of-the-art LLM4BS tasks (§4), comprising LLM4FUZZ [44], SMARTINV [29], BLOCKGPT [26]. Finally, we present an insightful discourse on the challenges presently faced within the ambit of LLM4BS, and proffer prospective trajectories for future research and development in this emergent field (§5).
@@
  • To the best of our knowledge, after a meticulous review of the existing literature, we conduct the first systematic examination focusing on the application of Large Language Models to tasks within the realm of blockchain security, offering a pioneering exploration of the interplay between advanced AI and distributed ledger systems.
  • In our comprehensive survey, we meticulously chronicle the current landscape of applications of Large Language Models in the domain of blockchain security. We delve into a detailed analysis of how Large Language Models are employed across various scenarios, from enhancing the reliability of smart contracts to fortifying the integrity of distributed ledger systems. This sheds light on the multifaceted contributions of this cutting-edge technology.
  • Based on our study, we rigorously compile and summarize a range of practical academic achievements related to the application of Large Language Models (LLMs) in strengthening blockchain security. We also propose several promising avenues for future research, anticipating that these will catalyze substantial advancements and innovations within this burgeoning intersection of fields.
@@ 2 Overview of LLM4BS
We provide basic knowledge about LLM4BS tasks in this section, including LLM applications in §2.1 and threats of blockchain security in §2.2.
@@ 2.1.1 Definition and Characteristics of LLMs
Furthermore, LLMs demonstrate an advanced understanding of the context within language, enabling them to generate responses or predictions that are sensitive to the surrounding textual context [66]. This contextual awareness is achieved through mechanisms such as attention mechanisms and positional encodings, which enable LLMs to attend to relevant parts of the input sequence and model long-range dependencies effectively [67].
@@ 2.1.2 Applications of LLMs in Various Domains
As depicted in Fig. 1, the versatility and efficacy of LLMs have led to their widespread adoption across diverse domains and applications, where they have demonstrated exceptional performance and utility [69]. Some notable applications of LLMs include:
@@ 2.2.2 Common Security Threats in Blockchain Systems
Auxiliary Service Vulnerabilities: Auxiliary services, such as wallets, exchanges, oracles, and decentralized applications (DApps), serve as entry points for attackers to exploit vulnerabilities and compromise the security of blockchain systems [116, 117]. Security breaches in auxiliary services, such as exchange hacks, wallet vulnerabilities, or oracle manipulation attacks, can lead to the loss of funds, unauthorized access to user data, or manipulation of on-chain transactions [118, 119]. Furthermore, the interconnected nature of auxiliary services within the blockchain ecosystem amplifies the impact of security breaches, as vulnerabilities in one service can propagate to others, resulting in widespread disruption and financial losses.
@@
3 Taxonomy of LLM4BS tasks
Figure 3: The applications of LLM on the task of blockchain security.
@@ 3.1 LLM as Code auditor on Smart Contracts
The application of LLM in the domain of smart contract code auditing and vulnerability detection can be succinctly encapsulated as follows: Advanced tools, such as SMARTINV [29], GPTScan [13], David et al. [30], Karanjai et al. [31], ContractArmor [32], Ortu et al. [33], ASSBert [34], PSCVFinder [35], LLM4Vuln [36], TrustLLMf [37], AuditGPT [38], Chen et al. [40], Jain et al. [41], PropertyGPT [39] and SolGPT [42]. As shown in Table.2, these tools powered by Large Language Models signify a monumental shift from traditional, pattern-based analysis methodologies towards more contextually aware and comprehensive inspection techniques. These cutting-edge tools extend their analytical prowess beyond static patterns by knitting together disparate threads of information, including the nuanced aspects of natural language documentation that detail the intended functions and transactional constructs of smart contracts.
The integration of code and contextual data through a multimodal lens equips such tools with the capacity to unravel complex logical oversights and identify subtle ”machine un-auditable” bugs, which would otherwise evade detection. By assimilating and interpreting the richer tapestry of human language explanations paired with code, LLM-based tools delve deeper into the intricate web of smart contract interactions. The profound understanding garnered from this approach not only sheds light on hidden vulnerabilities but also fortifies smart contracts against the myriad of risks that could lead to substantial financial repercussions.
In essence, the integration of Large Language Models in smart contract analysis marks a significant leap in safeguarding the infrastructural integrity of blockchain technology. It underscores an evolving landscape where artificial intelligence converges with software development practices to bolster security measures. This proactive identification and remediation of weaknesses within smart contracts, facilitated by the keen insights offered by LLMs, are instrumental in cementing trust and reliability in blockchain transactions—hence mitigating potential financial liabilities and reinforcing the bedrock of digital contracts.
@@
Chen et al. [40] Contract vulnerability detection ChatGPT etc. arXiv,2023
@@
PropertyGPT [39] Formal verification automation GPT-4 arXiv,2024
SolGPT [42] Contract vulnerability detection GPT-2 ICA3PP,2023
@@ 3.2 LLM as Analyzers for abnormal transaction
The application of LLMs for blockchain transaction analysis, such as BLOCKGPT [26] and Nicholls et al. [43], underscores their crucial role in conducting real-time monitoring to detect signs of irregular or suspicious behavior. These tools in Table 3 represent a significant advancement in the field, as they provide a more dynamic and adaptable approach to identifying potential threats within blockchain transactions.
@@
In addition to identifying potentially fraudulent transactions, LLMs also contribute to risk assessment and regulatory compliance. By analyzing the transaction data against current compliance standards and risk models, LLMs can assist financial institutions in managing their risk exposure and adhering to anti-money laundering (AML) and know-your-customer (KYC) regulations. Their sophisticated analysis capabilities can provide valuable insights to compliance officers and regulatory bodies, allowing for a more proactive approach to detecting and preventing financial crimes.
@@ 3.3 LLM as Fuzzer for Smart Contract
Large Language Models (LLMs) have been increasingly employed to elevate the process of fuzzing, particularly in the realm of smart contract security analysis, such as LLM4FUZZ [44] and ACFIX [45]. This methodology in Table 4 involves utilizing LLMs to accurately assess the complexity and vulnerability likelihood of specific code regions within a smart contract. Consequently, these metrics serve to guide the direction and focus of fuzzers, steering them toward code segments that are more likely to harbor potential security threats.
@@
Moreover, this refined fuzzing technique allows for the integration of user-defined invariants and manually inserted assertions to monitor and manage the state during fuzzing. This approach can reduce the exploration overhead and improve the detection of more profound logical issues that regular fuzzing routines might miss. Evaluations of this LLM-enhanced fuzzing method within real-world decentralized finance (DeFi) projects have demonstrated its effectiveness, outperforming baseline fuzzing parameters and uncovering significant vulnerabilities. These vulnerabilities, if left undetected and exploited, could potentially result in substantial financial losses.
@@ 3.4 LLM as Developer for Smart Contract
Storhaug et al. [27] Vulnerability-constrained decoding GPT-J-6B ISSRE,2023
@@
Petrovic et al. [49] Smart contract generation ChatGPT ICSEng,2023
Zhao et al. [50] AI programming assistant GPT-3.5 arXiv,2024
Haque et al. [51] Norm extraction ChatGPT arXiv,2024
Recent studies in Table 5, such as Storhaug et al. [27], karanjai et al. [31], Petrovic et al. [49], Zhao et al. [50], Haque et al. [51], MazzumaGPT [46], Du et al. [47] and GPTutor [48], have begun to scrutinize the efficacy and reliability of Large Language Models (LLMs) like ChatGPT and Google Palm2 in the automated generation of smart contracts. These smart contracts are integral to the blockchain ecosystem, executing agreements without the need for intermediaries, and their accuracy and security are paramount. The research primarily constructs a testing framework that assesses smart contracts on multiple fronts, i.e., validity, correctness, efficiency, security, and maintainability.
@@
These works point to the necessity for comprehensive analysis and improvement in the methodologies employed by LLMs. There is optimism that future iterations of LLMs, with better training and prompt design considerations, could enhance the quality and security of AI-generated smart contracts. It also hints at the potential for these tools to revolutionize contract generation by reducing the time and effort required, while flagging the urgent need for more robust security measures and testing methods.
@@ 3.5 LLM as Participants for Cryptocurrency community
Governance emerges as a major theme, as LLMs could contribute significantly to the structuring and transparency of this largely unregulated space. The first document outlines the broader governance challenges faced by AI systems, suggesting blockchain as a viable solution to introduce verifiability and accountability. On the other hand, the limitations of LLMs in capturing the complexities of legal reasoning are highlighted, a concern that is echoed across the three studies to varying degrees.
@@
In essence, the collective narrative from the three documents converges on the premise that LLMs hold transformative potential for the cryptocurrency community’s governance and legal sectors but must overcome challenges in understanding before they can be fully trusted in autonomous roles.